### v6

Abstract
Reinforcement learning has recently gained unprecedented popularity, yet it still grapples with sample inefficiency. Addressing this challenge, federated reinforcement learning (FedRL) has emerged, wherein agents collaboratively learn a single policy by aggregating local estimations. However, this aggregation step incurs significant communication costs. In this paper, we propose CompFedRL, a communication-efficient FedRL approach incorporating both periodic aggregation and (direct/error-feedback) compression mechanisms. Specifically, we consider compressed federated $Q$-learning with a generative model setup, where a central server learns an optimal $Q$-function by periodically aggregating compressed $Q$-estimates from local agents. For the first time, we characterize the impact of these two mechanisms (which have remained elusive) by providing a finite-time analysis of our algorithm, demonstrating strong convergence behaviors when utilizing either direct or error-feedback compression. Our bounds indicate improved solution accuracy concerning the number of agents and other federated hyperparameters while simultaneously reducing communication costs. To corroborate our theory, we also conduct in-depth numerical experiments to verify our findings, considering Top-$K$ and Sparsified-$K$ sparsification operators.
Keywords:
Federated Reinforcement Learning Communication Efficiency Direct Compression Error-feedback Compression.
1 Introduction
Reinforcement learning (RL) [56] has advanced many progresses in solving several decision-making applications, such as game playing [49], optimal control [10], autonomous driving [70], and recommendation systems [65]. In traditional RL algorithms, an agent learns and makes the best sequential actions within a Markov Decision Process (MDP) that maximize long-term rewards, even without the full knowledge of its model parameters. This learning process involves the agent repeatedly interacting with the environment, generating data that can then be used to refine its decision-making strategy.
However, training RL models under high-dimensional environments is challenging, because the RL models require (1) high volumes of storage being directly proportional to the size of state and action spaces, and (2) high sample complexity to find an optimal policy, i.e. a large number of interactions between the agent and the environment [11, 9]. Training RL models with limited data often leads to poor performance in high output variability, while sequentially generating vast volumes of data is computationally expensive. Consequently, many practical RL implementations resort to parallel data sampling from the environment by multiple agents [48].
Recently, many RL algorithms have been modified to leverage federated learning (FL), known as federated RL (FedRL), to obtain faster training performance and improved accuracy. FL is a popular multi-agent framework that ensures privacy and security [45] since it does not require the agents to transmit their private data over the network, which is computationally prohibitive, especially when the data is sizeable, and the network is resource-constrained. At each communication round of FedRL algorithms, the central server updates the model parameters, based on aggregations of local model updates transmitted by all the agents. Previous studies on FedRL algorithms primarily focus on addressing challenges such as robustness to adversarial attacks [69, 29], environment heterogeneity [34, 63, 71], behavior policy heterogeneity [68], and sample complexities [26, 37]. Unfortunately, none of these studies addresses the communication performance bottleneck of FedRL algorithms that interact with high-dimensional environments.
A commonly used approach to reduce communication costs is to apply compression, e.g., sparsification and quantization, alongside periodic aggregation. Compressing the information reduces its size before it is transmitted and is beneficial for running optimization algorithms over bandwidth-limited networks. The impact of compression on optimization algorithms under both centralized and federated settings has been extensively studied, e.g. in [1, 2, 64, 54, 36, 14]. Unfortunately, very limited works have studied compressed FedRL algorithms. Multi-agent RL with compressed updates has been studied in [47]. However, the result in [47] is limited to TD learning with linear function approximation, and they neither utilized periodic averaging nor analyzed direct compression.
Contributions.
  • CompFedRL. In Section 4, we propose CompFedRL, a novel FedRL algorithm that incorporates both periodic aggregation and (direct/error-feedback) compression mechanisms to guarantee strong convergence while mitigating substantial costs of communicating local $Q$-estimates.
  • Non-asymptotic convergence results for CompFedRL under UnbiasedComp and BiasedComp compression. We provide the convergence analysis of CompFedRL, characterizing the impact of these mechanisms on convergence behaviors. Through a finite-time analysis of our algorithm, we demonstrate robust convergence behaviors when employing either direct or error-feedback compression techniques in Section 5. Our theoretical bounds highlight the enhanced solution accuracy achieved by CompFedRL, particularly concerning the number of agents and other federated hyperparameters, while simultaneously reducing communication costs.
  • Empirical validation. To validate our theoretical observations, in Section 6, we conduct extensive numerical experiments, leveraging sparsification operators such as Top-$K$ and Sparsified-$K$. These experiments corroborate the effectiveness of CompFedRL in practice, further reinforcing the significance of our contributions in advancing communication-efficient FedRL algorithms. Our experiments’ code is publicly available in the following link: https://github.com/AliBeikmohammadi/CompFedRL.
2 Related Work
In what follows, we delve into the most pertinent threads of related work, with a particular focus on theoretical contributions in the field. We present prior works closely related to our paper in RL and communication-efficient learning.
2.1 Single-agent RL Algorithms
Existing works have extensively investigated the convergence guarantees of many RL algorithms, such as TD-learning and $Q$-learning algorithms, primarily on the single-agent environment under various conditions. TD-learning algorithms are shown to have asymptotic convergence in the on-policy setting by [59, 58, 16] and in the off-policy setting by [42, 72], while the algorithms attain non-asymptotic convergence (or finite-sample bounds) in the on-policy setting by [24, 38, 15, 53, 32, 19] and in the off-policy setting by [18, 19]. Furthermore, initially proposed by [66], the $Q$-learning algorithm has been extensively studied its convergence properties in varied settings. For instance, on the one hand, the asymptotic convergence of the $Q$-learning algorithm was established by [60, 33, 13, 57, 17, 16]. On the other hand, the finite-time convergence of $Q$-learning algorithm under synchronous sampling, also known as RL with a generative model, was proved by [28, 5, 62, 18, 39, 30], and under asynchronous sampling (known as Markovian sampling) was studied by [28, 5, 51, 39, 40, 19]. Our work operates within RL with a generative model, but focuses on a multi-agent collaborative federated setup, emphasizing communication-efficient methodologies, which requires an analysis framework different from traditional single-agent RL analyses.
2.2 Distributed and Federated RL Algorithms
Several recent studies have introduced distributed variants of RL algorithms aimed at expediting training processes [48, 27, 3]. Theoretical examinations of the convergence properties of these distributed RL algorithms have also been explored in recent literature across various scenarios, including decentralized stochastic approximation [25, 55, 61, 73, 47], TD learning with linear function approximation [63], and off-policy TD in actor-critic algorithms [21]. Additionally, [20, 52] have conducted analyses on the finite-time convergence of distributed actor-critic algorithms. Nevertheless, all of these investigations focus on decentralized setups, where agents communicate with their neighbors following each local update. On the other hand, there are some studies that consider a federated setting, with each agent performing multiple local updates between successive communication rounds, thereby resulting in communication savings. This includes TD learning with linear function approximation [23, 22, 37], off-policy TD learning [37], SARSA with linear function approximation [71], and $Q$-learning [37, 68, 34]. Our work, akin to the setup of [68], centers on a federated $Q$-learning algorithm but sets itself apart by integrating various compression strategies to enhance communication efficiency.
2.3 Communication-Efficient Learning Algorithms
Many compression schemes, i.e. sparsification and/or quantization, have been used for optimization algorithms to train state-of-the-art learning models (e.g. AlexNet, VGG19, and ResNet152) under resource-constrained networks [1, 67, 64, 44]. These compression schemes can be either unbiased (such as QSGD [1], TernGrad [67], and ATOMO [64]) or biased (e.g. Top-$K$[2, 7], normalized gradient [43] and sign compression [12]). Despite communication benefits, optimization algorithms with compression often have poor convergence performance. For instance, the biased compression algorithms are shown to fail to converge for simple problems [14, 36]. To address the problem of direct compression, several studies [54, 36, 2, 35, 41, 31, 46, 8] have proposed error-feedback mechanisms which leverage memory of previous compression errors to mitigate compression biases for distributed and federated learning algorithms under various settings. None of the mentioned works has investigated communication-efficient learning in the context of RL. The only existing study, [47], examines error-feedback compression, focusing on TD learning with linear function approximation in distributed learning, contrasting with our synchronous federated $Q$-learning approach, where aggregation occurs periodically. We investigate both direct and error-feedback compression operators, offering a thorough FedRL convergence analysis.
2.4 Notation
We denote by $\mathbb{N}$, $\mathbb{N}_{0}$, and $\mathbb{R}$ the sets of natural numbers, natural numbers including zero, and real numbers, respectively. The set $\{a,a+1,...,b\}$ is represented as $[a,b]$, where $a,b\in\mathbb{N}_{0}$ and $a\leq b$. For a vector $v\in\mathbb{R}^{d}$, $\|v\|_{2}$ and $\|v\|_{\infty}$ are its $\ell_{2}$-norm and $\ell_{\infty}$-norm, respectively.
3 Preliminaries and Background
Within this section, we present the mathematical framework and fundamental concepts of MDPs.
3.1 Discounted Infinite-horizon MDP
We consider a discounted infinite-horizon MDP, denoted by a quintuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}=\{1,2,...,|\mathcal{S}|\}$ and $\mathcal{A}=\{1,2,...,|\mathcal{A}|\}$ represent the finite state space and action space, respectively. The transition probability kernel $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ determines the likelihood $P\left(s^{\prime}\mid s,a\right)$ of transitioning from state $s$ to state $s^{\prime}$ upon taking action $a$. The reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ assigns an immediate reward $r(s,a)$ for executing action $a$ in state $s$. Additionally, the discount factor $\gamma\in(0,1)$ adjusts the weighting of future rewards.
3.2 Policy, and $Q$-function
A policy $\pi$ is a mapping $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$, where $\pi(a\mid s)$ denotes the probability of selecting action $a$ given state $s$. Given a policy $\pi$, the value function $V^{\pi}:\mathcal{S}\rightarrow\mathbb{R}$ quantifies the expected discounted cumulative reward starting from an initial state $s$ and is defined as
$V^{\pi}(s)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r\left(s_{k},a_{k}\right)\mid s_{0}=s\right],\forall\leavevmode\nobreak\ s\in\mathcal{S}.$ (1)
Here, the expectation is computed over the randomness of the trajectory $\left\{s_{k},a_{k},r_{k}\right\}_{k=0}^{\infty}$, sampled according to the transition kernel (i.e., $)s_{k+1}\sim P\left(\cdot\mid s_{k},a_{k}\right)$ and policy $\pi$ (i.e., $)a_{k}\sim\pi\left(\cdot\mid s_{k}\right)$ for any $k\geq 0$. Similarly, the state-action value function, denoted as the $Q$-function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, measures the expected discounted cumulative reward from an initial state-action pair $(s,a)$:
$Q^{\pi}(s,a)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r\left(s_{k},a_{k}\right)\mid s_{0}=s,a_{0}=a\right],\forall\leavevmode\nobreak\ (s,a)\in\mathcal{S}\times\mathcal{A}.$ (2)
3.3 Optimal Policy and Bellman Operator
An optimal policy $\pi^{\star}$ maximizes the value function uniformly across all states. Note that the existence of such an optimal policy is always assured [50, 56], and it simultaneously maximizes the $Q$-function. The corresponding optimal value function and $Q$-function are represented as $V^{\star}=V^{\pi^{\star}}$ and $Q^{\star}=Q^{\pi^{\star}}$, respectively. It is a well-established fact that the optimal $Q$-function $Q^{\star}$ can be determined as the unique fixed point of the population Bellman operator $\mathcal{T}:\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$, which is defined as:
$\mathcal{T}(Q)(s,a)=r(s,a)+\gamma\underset{s^{\prime}\sim P(\cdot\mid s,a)}{\mathbb{E}}\left[\max_{a^{\prime}\in\mathcal{A}}Q\left(s^{\prime},a^{\prime}\right)\right].$ (3)
3.4 RL with a Generative Model
In the context of learning, the transition kernel $P$ is typically unknown. Thus computing the Bellman operator precisely is infeasible. Instead, we assume access to a simulation engine that generates samples. In this paper, we focus on the synchronous or generative setting, where at each time step $k$ and for each state-action pair $(s,a)$, a sample $s_{k}(s,a)$ is observed, drawn according to the transition function $P(\cdot\mid s,a)$ (i.e., $s_{k}(s,a)\sim P(\cdot\mid s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$). Also, convergence guarantees for the synchronous setting can be easily translated for the asynchronous setting, e.g. by using cover times of Markov chains [28, 4].
In the synchronous form of the $Q$-learning algorithm, a sequence of $Q$-tables $\left\{Q_{k}\right\}_{k\geq 1}$ is generated according to:
$Q_{k+1}=\left(1-\eta\right)Q_{k}+\eta\mathcal{T}_{k}\left(Q_{k}\right),$ (4)
where $\eta\in(0,1]$ is a learning rate, and the operator $\mathcal{T}_{k}:\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ is an empirical Bellman operator with its $(s,a)$-entry evaluated by:
$\mathcal{T}_{k}(Q)(s,a)=r(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q\left(s_{k},a^{\prime}\right),$ (5)
where $s_{k}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ is a random matrix indexed by state-action pairs $(s,a)$, with each entry $s_{k}(s,a)$ drawn from the probability distribution $P(\cdot\mid s,a)$. Note that $\mathbb{E}\left[\mathcal{T}_{k}(Q)\right]=\mathcal{T}(Q)$ for any fixed $Q$, indicating that the empirical Bellman operator (5) is an unbiased estimate of the population Bellman operator (3).
4 CompFedRL
Algorithm 1 Compressed FedRL with a Generative Model (CompFedRL)
1: inputs: Learning rate $\eta\in(0,1]$, Discount factor $\gamma\in(0,1)$, Federated parameter $\beta\in(0,1]$, Number of agents $I$, Number of local epochs $K$, Number of communication rounds $T$, Compress operator $\in\{{\textsf{UnbiasedComp}},{\textsf{BiasedComp}}\}$.
2: initialization: Server broadcasts $\overline{Q}_{0}=Q_{0}$.
3: if Compress operator is BiasedComp then
4:      Initial errors $e_{0}^{(i)}=0,\forall i\in[I]$.
5:end if
6: for each round $t=0,\ldots,T-1$ do
7:      for each agent $i\in[I]$ in parallel do
8:          Agent $i$ initializes $Q_{t,0}^{(i)}=\overline{Q}_{t}$
9:          for $k=0,\ldots,K-1$ do
10:               Draw $s_{t,k+1}^{(i)}\left(s,a\right)\sim{P}{\left(.|s,a\right)}$ for all ${\left(s,a\right)}\in\mathcal{S}\times\mathcal{A}$.
11:               Update local $Q_{t,k+1}^{(i)}$ according to (6)
12:         end for
13:          if Compress operator is UnbiasedComp then
14:               Compute $h_{t}^{(i)}=\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t})$ and send back to the server.
15:         end if
16:          if Compress operator is BiasedComp then
17:               Compute $h_{t}^{(i)}=\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t}+e_{t}^{(i)})$ and send back to the server.
18:               Update $e_{t+1}^{(i)}=Q_{t,K}^{(i)}-\overline{Q}_{t}+e_{t}^{(i)}-h_{t}^{(i)}$.
19:         end if
20:     end for
21:      Server computes and broadcasts global $\overline{Q}_{t+1}=\overline{Q}_{t}+\frac{\beta}{I}\sum_{i=1}^{I}{h_{t}^{(i)}}$
22:end for
We introduce communication-efficient FedRL with a generative model called CompFedRL. This algorithm utilizes periodic aggregation and compression mechanisms to obtain provably strong convergence guarantees while reducing communication bandwidth for transmitting local $Q$-estimates from the agents to the central server. At each communication round $t=[0,T-1]$ of the algorithm, every agent $i\in[1,I]$ receives a global $Q$-table $\overline{Q}_{t}$ from the central server and initializes its $Q$-estimate table $Q^{(i)}_{t,0}=\overline{Q}_{t}$. Each agent in parallel updates all entries of its $Q$-estimates for $K$ local epochs according to:
$Q_{t,k+1}^{(i)}\left(s,a\right)=\left(1-\eta\right)Q_{t,k}^{(i)}\left(s,a\right)+\eta\mathcal{T}^{(i)}_{k}(Q_{t,k}^{(i)})\left(s,a\right)$ (6)
where $\eta$ is a positive learning rate, and $\mathcal{T}^{(i)}_{k}$ is the empirical Bellman operator at local epoch $k$ defined by
$\mathcal{T}^{(i)}_{k}(Q)\left(s,a\right)=r\left(s,a\right)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q\left(s_{t,k+1}^{(i)}\left(s,a\right),a^{\prime}\right),\quad s_{t,k+1}^{(i)}\left(s,a\right)\sim P(\cdot|s,a).$ (7)
After running for $K$ local epochs, each agent transmits to the server the compressed progress between the local $Q$-estimate and the global $Q$-table denoted by $h_{t}^{(i)}$. Here, $h_{t}^{(i)}=\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t})$ if $\textsf{Compress}(\cdot)$ is UnbiasedComp, and $h_{t}^{(i)}=\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t}+e_{t}^{(i)})$ if $\textsf{Compress}(\cdot)$ is BiasedComp where $e_{t}^{(i)}$ maintains the memory of compression errors defined by $e_{t+1}^{(i)}=Q_{t,K}^{(i)}-\overline{Q}_{t}+e_{t}^{(i)}-h_{t}^{(i)}$.
Finally, after receiving all $h_{t}^{(i)}$ from every agent, the central server updates the global $Q$-table $\overline{Q}_{t+1}=\overline{Q}_{t}+({\beta}/{I})\sum_{i=1}^{I}h_{t}^{(i)}$, where $\beta\in[0,1]$ is the federated parameter. The full description of CompFedRL  is in Algorithm 1. Further note that Algorithm 1 becomes QAvg [34] and FedSyncQ [68] when $\textsf{Compress}(v)=v$ and $\beta=1$, and recovers a centralized synchronous $Q$-learning algorithm analyzed by [39, 5, 18, 28] by choosing $\textsf{Compress}(v)=v$, $\beta=1$ and $I=1$.
4.1 Compression Options for CompFedRL
The choices of compression operators deployed in CompFedRL are flexible. We define two types of compression operators that are useful for our analysis.
Definition 1 (UnbiasedComp)
$\textsf{Compress}(v):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is UnbiasedComp if it satisfies the following: for $j\in[1,d]$ and $v\in\mathbb{R}^{d}$ (1) $\mathbf{E}[\textsf{Compress}(v)]=v$, (2) $\mathbf{E}|\textsf{Compress}(v)_{j}-v_{j}|^{2}\leq q_{2}|v_{j}|^{2}$ for some $q_{2}>0$, and (3) $\|\textsf{Compress}(v)-v\|_{\infty}\leq q_{\infty}\|v\|_{\infty}$ for some $q_{\infty}>0$.
Sparsified-$K$ is one popular operator satisfying Definition 1. The Sparsified-$K$ operator transmits on average $K$ elements of a vector $v\in\mathbb{R}^{d}$, which is defined by
$[\textsf{Compress}(v)]_{j}=(v_{j}/p_{j})\xi_{j},$ (8)
where $v_{j}$ is the $j^{\text{th}}$-element of $v\in\mathbb{R}^{d}$, $\xi_{j}\sim\texttt{Bernouli}(p_{j})$ and $K=\sum_{j=1}^{d}p_{j}$. The Sparsified-$K$ operator captures several variants of sparsification in communication-efficient training. For example, Sparsified-$K$ with $p_{j}=|v_{j}|/\|v\|_{q}$ and with $q=2$, $q=\infty$ and $q\in(0,\infty)$ is 1-bit QSGD in [1], Terngrad in [67] and $\ell_{q}$-quantization in [64]. Finally, the next lemma shows that Sparsified-$K$ follows Definition 1.
Lemma 1
Let $\textsf{Compress}(v)$ be Sparsified-$K$ and $p_{\min}=\min_{j=[1,d]}p_{j}$. Then, $\textsf{Compress}(v)$ is UnbiasedComp with $q_{2}=1/p_{\min}-1$ and $q_{\infty}=\max(1/p_{\min}-1,1)$.
Next, we provide the definition of BiasedComp, which captures biased compression operators of interest.
Definition 2 (BiasedComp)
$\textsf{Compress}(v):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is BiasedComp if it satisfies $\|\textsf{Compress}(v)-v\|_{\infty}\leq(1-\alpha)\|v\|_{\infty}$ for some $\alpha\in(0,1]$ and for all $v\in\mathbb{R}^{d}$.
This definition requires the contraction property in the $\ell_{\infty}$-norm, rather than the $\ell_{2}$-norm in existing works, e.g. [2, 36, 54, 47]. One example that satisfies Definition 2 is Top-$K$. Instead of transmitting on average $K$ elements, Top-$K$ sends $K$ elements with the highest absolute magnitude from a vector $v\in\mathbb{R}^{d}$. This can be done by applying the following operator
$[\textsf{Compress}(v)]_{i}=\begin{cases}v_{i}&\text{if }i\in I_{K}(v)\\ 0&\text{otherwise}\end{cases}$ (9)
Here, $I_{K}(v)$ is the index set for $K$ largest components of $v$ in the absolute magnitude. Finally, the Top-$K$ operator achieves the contraction in the $\ell_{\infty}$-norm in the next lemma.
Lemma 2
Let $\textsf{Compress}(v)$ be Top-$K$ and $\max_{j\in I_{K}^{c}(v)}|v_{j}|<\|v\|_{\infty}$ where $I_{K}^{c}(v)=[1,d]-I_{K}(v)$. Then, $\textsf{Compress}(v)$ is BiasedComp with $\alpha=1-{\max_{j\in I_{K}^{c}(v)}|v_{j}|}/{\|v\|_{\infty}}$.
5 Convergence Analysis
To this end, we provide the non-asymptotic convergence of CompFedRL with UnbiasedComp and BiasedComp.
5.1 CompFedRL with UnbiasedComp
The first theorem presents the convergence for CompFedRL with UnbiasedComp.
Theorem 5.1
Consider Algorithm 1 with UnbiasedComp. Let $\beta,\eta\in(0,1]$ and $0\leq\|Q_{0}\|_{\infty}\leq\frac{1}{1-\gamma}$. Then, with probability at least $1-\delta$
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $\rho^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\sqrt{\frac{\eta}{I}}e_{1}+\frac{2\gamma}{C}+\frac{1}{\sqrt{I}}e_{2},$
where $\rho=1-\beta+\beta(1-\eta)^{K}$, $e_{1}=\frac{4\gamma}{C}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\right)$, $e_{2}=\frac{1}{(1-(1-\eta)^{K})}\left(\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right)$, and $C=(1-\gamma)(1-(1-\eta)^{K})$.
Discussions on Theorem 5.1:
Theorem 5.1 establishes non-asymptotic high-probability convergence with the linear rate depending on federated parameters $\beta,\eta,K$ and three residual error terms for CompFedRL with UnbiasedComp. We make the following observations. First, there is a trade-off between convergence speed and solution accuracy when we adjust the learning rate $\eta$. In particular, increasing $\eta$ improves the convergence speed while increasing the first residual error term. Second, the first and last error term depends on $1/\sqrt{I}$, thus implying that CompFedRL with UnbiasedComp attains the speed-up with respect to the number of agents like multi-agent TD(0) learning in [47]. Third, the second error term is insignificant when the discount factor $\gamma$ is small. Fourth, Theorem 5.1 with $q_{2}=q_{\infty}=0$ provides the convergence for full-precision federated $Q$-learning.
We obtain the convergence for CompFedRL with UnbiasedComp and $K=1$ in the next corollary:
Corollary 1
Consider Algorithm 1 with UnbiasedComp and $K=1$ under the same settings as Theorem 5.1. Then, with probability at least $1-\delta$
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta\eta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\sqrt{\frac{1}{\eta I}}\bar{e}_{1}+\frac{2\gamma}{\eta(1-\gamma)}+\frac{1}{\eta\sqrt{I}}\bar{e}_{2},$
where $\bar{e}_{1}=\frac{4\gamma}{(1-\gamma)}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|T}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|T}{\delta}}\right)$, and
$\bar{e}_{2}=\left(\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right)$.
From Corollary 1, we obtain the lowest convergence bound by letting $\eta=1$, which yields:
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\sqrt{\frac{1}{I}}\bar{e}_{1}+\frac{2\gamma}{(1-\gamma)}+\frac{1}{\sqrt{I}}\bar{e}_{2}.$
From this bound, CompFedRL with $K=1$ and UnbiasedComp close to the identity operator (i.e. $q_{2}\rightarrow 0$ and $q_{\infty}\rightarrow 0$) and with $I\geq\max\left(B_{1},B_{2}\right)$ where $B_{1}=\left(\frac{4\gamma}{1-\gamma}\right)^{2}\frac{16\log(4|\mathcal{S}||\mathcal{A}|T/\delta)}{\epsilon^{2}}$ and $B_{2}=\left(\frac{4\gamma}{1-\gamma}\right)\frac{4\sqrt{\log(4|\mathcal{S}||\mathcal{A}|T/\delta)}}{\epsilon}$ achieves $\|\overline{Q}_{T}-Q^{\star}\|_{\infty}\leq\epsilon$ within $T=\frac{1}{\beta}\log(4\|\overline{Q}_{0}-Q^{\star}\|_{\infty}/\epsilon)$ iterations, when $\gamma\leq\frac{\epsilon}{8+\epsilon}$. If $\beta=(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})/\log(4\|\overline{Q}_{0}-Q^{\star}\|_{\infty}/\epsilon)$, then we have the $\mathcal{O}\left(\frac{1}{(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})}\right)$ complexity that is close to the lower bound for a single-agent $Q$-learning by [39].
5.2 CompFedRL with BiasedComp
The next theorem shows the convergence for CompFedRL with BiasedComp.
Theorem 5.2
Consider Algorithm 1 with BiasedComp. Let $\beta,\eta\in(0,1]$ and $0\leq\|Q_{0}\|_{\infty}\leq\frac{1}{1-\gamma}$. Then, with probability at least $1-\delta$
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $\rho^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{C}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}+\frac{2\gamma}{C}+\frac{2\beta(1-\alpha)}{\alpha(1-\gamma)}D,$
where $\rho=1-\beta+\beta(1-\eta)^{K}$, $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}$, $C=(1-\gamma)(1-(1-\eta)^{K})$, and $D=1+\frac{1+(1-\eta)^{K}}{1-(1-\eta)^{K}}$.
Discussions on Theorem 5.2:
From Theorem 5.2 CompFedRL with BiasedComp converges linearly with high probability with three residual error terms. The following observations from this theorem are similar to those from Theorem 5.1. First, there is a trade-off between convergence speed and solution accuracy of the algorithm when we adjust the learning rate $\eta$, the federated parameter $\beta$, and the number of local epochs $K$. In particular, the algorithm with $K\gg 1$ and $\beta,\eta\rightarrow 1$ achieves high speed at the price of low solution accuracy. Second, due to the dependency of the first error term in the convergence bound on $1/\sqrt{I}$, CompFedRL with BiasedComp achieves the speedup with respect to the number of agents similarly to CompFedRL with UnbiasedComp in Theorem 5.1 and to multi-agent TD(0) learning using error-feedback analyzed by [47]. Third, the last error term is due to the compression $\alpha$ and can be diminished by decreasing $\beta$ and increasing $\eta,K$, whereas the second error term is negligible when the discount factor $\gamma$ is close to zero. Fourth, we obtain the convergence for full-precision synchronous $Q$-learning when we let $\alpha=1$ in Theorem 5.2.
The next corollary presents the convergence of CompFedRL with BiasedComp when we let the number of local epoch $K=1$:
Corollary 2
Consider Algorithm 1 with BiasedComp and $K=1$ under the same settings as Theorem 5.2. Then, with probability at least $1-\delta$
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta\eta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{1-\gamma}\sqrt{\frac{1}{\eta I}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}e_{1}+\frac{2\gamma}{\eta(1-\gamma)}.$
From Corollary 2, CompFedRL with $K=1$ and BiasedComp, unlike UnbiasedComp, does not have the residual error term due to compression. Furthermore, CompFedRL with $K=1$ and BiasedComp  attains the lowest convergence bound when we let $\eta=1$, which yields:
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{1-\gamma}\sqrt{\frac{1}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}e_{1}+\frac{2\gamma}{1-\gamma}.$
In conclusion, this algorithm reaches $\|\overline{Q}_{T}-Q^{\star}\|_{\infty}\leq\epsilon$ within $T=\frac{1}{\beta}\log(3\|\overline{Q}_{0}-Q^{\star}\|/\epsilon)$ iterations when $I=\frac{16}{(1-\gamma)^{2}}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}$ and $\gamma\leq\frac{\epsilon}{6+\epsilon}$. If $\beta=(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})/\log(3\|\overline{Q}_{0}-Q^{\star}\|/\epsilon)$, then we achieve the $\mathcal{O}\left(\frac{1}{(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})}\right)$ complexity which nearly matches the lower bound for the single-agent $Q$-learning by [39].
6 Experiments
In this section, we carry out numerical experiments to showcase the effectiveness of the CompFedRL algorithm.
6.1 Setup
We examine five variations of the noisy grid-world environment, some of which contain walls (as illustrated in Figure 5 of Appendix). Details regarding these environments and more results can be found in Appendix 0.C.1 and Appendix 0.C.2, respectively. The $Q$-function is initialized with zero entries, and we set $\gamma=0.8$. After each round of communication $t$, we evaluate the estimated $Q$ through the root mean square error (RMSE) between the estimated $\overline{Q}_{t}$ and the optimal $Q^{*}$.
To measure the communication efficiency of compression, we not only monitor the RMSE in communication rounds but also track the number of bits used by each agent. We evaluate communicated bits by considering floating point precision (FPP) to be 32 bits. Then, in each round of communication, for FedRL (i.e., CompFedRL without compression), CompFedRL with Top-$K$, and CompFedRL with Sparsified-$K$, the payload consumes $|S||A|\times FPP$ bits, $K\times\left(\lceil\log_{2}(|S||A|)\rceil+FPP\right)$ bits, and $K^{\prime}\times\left(\lceil\log_{2}(|S||A|)\rceil+FPP\right)$ bits, respectively. Here, $K^{\prime}$ is the size of the selected set by Sparsified-$K$ to be sent, and the factor $\lceil\log_{2}(|S||A|)\rceil$ comes from indicating $K$ ($K^{\prime}$) indices in the dimension of the state-action space.
Figure 1: Impact of compression: RMSE of the $Q$-estimates for both FedRL [68, 34] (without compression) and CompFedRL to solve Map11×$11\times 1$ 11 task. Here, $I=50$, $\eta=0.01$, $\beta=0.8$, and (Left) $K=1$; (Right) $K=10$.
Figure 2: Impact of the number of agents and local epochs: RMSE of the $Q$-estimates for CompFedRL under Top-50 sparsification to solve Map11×$11\times 1$ 11 task. Here, (Left) $T\times K=10000$, $\eta=0.01$, and $\beta=1$; (Right) $T=10000$, $K=1$, and $\beta=0.8$.
Figure 3: Impact of federated parameter: RMSE of the $Q$-estimates for CompFedRL under (Left) Top-50  and (Right) Sparsified-50  sparsification to solve Map11×$11\times 1$ 11 task. Here, $I=50$, $T=10000$, $K=1$, and $\eta=0.1$.
Figure 4: Impact of learning rate: RMSE of the $Q$-estimates for CompFedRL under (Left) Top-50  and (Right) Sparsified-50  sparsification to solve Map11×$11\times 1$ 11 task. Here, $I=50$, $T=10000$, $K=1$, and $\beta=0.8$.
6.2 Communication Efficiency of CompFedRL
Figure 1 demonstrates the impact of both Sparsified-$K$ and Top-$K$ on reducing communication payload with the server and on convergence. Transitioning from local epochs of $K=1$ to $K=10$ showcases consistent convergence behavior despite a tenfold reduction in communication frequency, underscoring the practicality of both periodic aggregation and compression mechanisms. A comparative analysis between Sparsified-$K$ and Top-$K$ reveals that Top-$K$ demonstrates lower residual error and converges to a point comparable to FedRL, as discussed in Section 5.2, where it was highlighted that BiasedComp does not incur a residual error term due to compression. Conversely, with Sparsified-$K$, an increase in compression leads to a corresponding increase in residual error.
6.3 Convergence Speedup
Local updating and the number of agents improve the performance of CompFedRL, as shown in Figure 2. The left plot shows substantial communication reduction by local updating (e.g. $10$ times bits saving for $K=10$ against $K=1$) and confirms Theorem 5.1 and 5.2 that, given a fixed step size, an increase in the number of agents lowers residual error, improving convergence performance in RMSE. From the right plot, an interesting observation is that a higher number of agents allows for larger learning rates, which accelerates the convergence speed to attain almost the same target $\epsilon$-accuracy.
6.4 Impact of Federated Parameter $\beta$ and Learning Rate $\eta$
Figure 3 and 4 demonstrate the trade-off between the convergence speed and solution accuracy when adjusting $\eta$ and $\beta$. Increasing $\eta$ improves the speed while worsening the accuracy, as confirmed by Theorem 5.1 and 5.2 (see Figure 4). Not captured by our theory, adjusting $\beta$ shows the same trade-off as $\eta$ as shown by Figure 3. Changes in $\eta$ affect residual error higher than adjustments in $\beta$, as supported by our non-asymptotic results emphasizing the dependency of most error terms on $\eta$. Finally, thanks to its error-feedback mechanism, Top-$K$ shows greater robustness than Sparsified-$K$ to these parameter tunings.
7 Conclusion
We have proposed CompFedRL, a communication-efficient federated $Q$-learning algorithm that utilizes local epochs and compression to attain strong convergence performance while saving communication costs. Under synchronous settings, we derive the non-asymptotic convergence of the algorithm using both unbiased and biased compression. Our theory suggests speed up with respect to the number of agents, and under certain choices of hyper-parameters CompFedRL attains the sample complexity that nearly matches the lower bound by [39]. Experiments on the noisy grid-world environment corroborate our analyses and illustrate the benefits of periodically compressed updating and the number of agents on convergence performance of CompFedRL in terms of communication efficiency and solution accuracy. Our empirical results also affirm the speed-and-accuracy trade-off when adjusting the learning rates and federated parameters.
Acknowlegments
This work was partially supported by the Swedish Research Council through grant agreement no. 2020-03607 and in part by Sweden’s Innovation Agency (Vinnova). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers Centre for Computational Science and Engineering (C3SE) partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Supplementary Material
For detailed appendices and associated code, please refer to the full version of this paper available on Springer Link and arXiv [6], as well as the code repository on GitHub: https://github.com/AliBeikmohammadi/CompFedRL.
Appendix 0.A Proof of Main Results
We present proofs for main lemmas and theorems in this section.
0.A.1 Useful lemmas
We begin by stating useful lemmas for our analysis. The first lemma proves the convergence bound for a non-negative sequence $V_{k}$ that satisfies $V_{k+1}\leq\rho V_{k}+e$ for $e>0$ and $\rho\in(0,1)$.
Lemma 3
Let the non-negative sequence $\{V_{k}\}_{k\geq 0}$ satisfy $V_{k+1}\leq\rho V_{k}+e$. If $\rho\in(0,1)$ and $e>0$, then $V_{k}\leq\rho^{k}V_{0}+e/(1-\rho)$.
The next lemma provides high-probability bounds using Freeman’s inequality, which extends from Bernstein’s inequality.
Lemma 4
Let $E=\frac{\eta}{I}\sum_{t=0}^{T-1}(1-\eta)^{T-1-t}\sum_{i=1}^{I}w^{(i)}_{t}$ where each $w^{(i)}_{t}\in\mathbb{R}$ is independent of each other, and satisfy $\mathbf{E}w^{(i)}_{t}=0$, $\|w^{(i)}_{t}\|_{\infty}=W$ and $\mathbf{E}|w^{(i)}_{t}|^{2}\leq\sigma^{2}$. Then,
$\left|E\right|\leq\sqrt{\frac{8\eta\sigma^{2}}{I}\log\frac{2}{\delta}}+\frac{4W}{3I}\log\frac{2}{\delta}$
with probability at least $1-\delta$.
Proof
Define $z^{(i)}_{t}=\eta\sum_{t=0}^{T-1}(1-\eta)^{T-1-t}w_{t}^{(i)}$. Then, $E=(1/I)\sum_{i=1}^{I}z^{(i)}_{t}$ which satisfies the following properties
  1. 1.
    $\|z^{(i)}_{t}\|_{\infty}\leq\|w^{(i)}_{t}\|_{\infty}\leq W$, where the first inequality derives from the fact that $\sum_{t=0}^{T-1}(1-\eta)^{T-1-t}\leq\frac{1}{\eta}$. Note that this step is different from the step for proving (82) in [68], which is incorrect because they must bound $B_{t}(s,a)=\max|\sum_{k=1}^{K}z_{i}^{k}(s,a)|$, not $B_{t}(s,a)=\max|z_{i}^{k}(s,a)|$.
  2. 2.
    $\mathbf{E}[z^{(i)}_{t}]=0$, which comes from the fact that $\mathbf{E}w^{(i)}_{t}=0$.
  3. 3.
    Next, we have
    $W_{I}$ $=$ $\sum_{i=1}^{I}\mathbf{E}(z_{t}^{(i)})^{2}$
    $=$ $\sum_{i=1}^{I}\sum_{t=0}^{T-1}\eta^{2}(1-\eta)^{2(T-1-t)}\mathbf{E}(w_{t}^{(i)})^{2}$
    $\leq$ $I\sigma^{2}\sum_{t=0}^{T-1}\eta^{2}(1-\eta)^{2(T-1-t)}\leq\eta I\sigma^{2}.$
By Theorem 4 of [68] with $m=1$, we have
$\left|\sum_{i=1}^{I}z_{t}^{(i)}\right|\leq\sqrt{8\eta I\sigma^{2}\log\frac{2}{\delta}}+\frac{4W}{3}\log\frac{2}{\delta}$
with probability at least $1-\delta$. This implies that with probability at least $1-\delta$,
$\left|E\right|\leq\sqrt{\frac{8\eta\sigma^{2}}{I}\log\frac{2}{\delta}}+\frac{4W}{3I}\log\frac{2}{\delta}.$
By using Lemma 4, we obtain a high-probability bound for $\left\|\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right\|_{\infty}$ for Algorithm 1 with both UnbiasedComp and BiasedComp as follows:
Lemma 5
Consider Algorithm 1. Then, with probability at least $1-\delta/T$
$\left\|\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right\|_{\infty}$$\leq$ $(1-\eta)^{K}\|\overline{Q}_{t}-Q^{\star}\|_{\infty}+\frac{4\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}+\frac{2\gamma}{1-\gamma},$
where $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}$.
Proof
Define $\Delta_{t,k}=\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,k}-Q^{\star}$. Then,
$\Delta_{t,k+1}=(1-\eta)\Delta_{t,k}+\frac{\eta}{I}\sum_{i=1}^{I}[r+\gamma P^{(i)}_{t,k+1}V^{(i)}_{t,k}-Q^{\star}],$
where the local transition matrix $P_{t,k+1}^{(i)}$ is $1$ if $s^{\prime}=s^{(i)}_{t,k+1}(s,a)$ and 00 otherwise, and $V_{t,k}^{(i)}(s)=\max_{a^{\prime}\in\mathcal{A}}Q^{(i)}_{t,k}(s,a^{\prime})$. Since $Q^{\star}=r+\gamma PV^{\star}$,
$\Delta_{t,k+1}=(1-\eta)\Delta_{t,k}+\frac{\gamma\eta}{I}\sum_{i=1}^{I}(P_{t,k+1}^{(i)}-P)V^{(i)}_{t,k}+\frac{\gamma\eta}{I}\sum_{i=1}^{I}P(V^{(i)}_{t,k}-V^{\star}).$
From the definition of the $\ell_{\infty}$-norm and by the triangle inequality,
$\|\Delta_{t,K}\|_{\infty}$$\leq$ $(1-\eta)^{K}\|\Delta_{t,0}\|_{\infty}+\|E^{1}_{t}\|_{\infty}+\|E^{2}_{t}\|_{\infty},$
where $E^{1}_{t}=\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}(P_{t,k+1}^{(i)}-P)V^{(i)}_{t,k}$ and $E^{2}_{t}=\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}P(V^{(i)}_{t,k}-V^{\star})$.
To complete the proof, we must bound $\|E^{1}_{t}\|_{\infty}$ and $\|E^{2}_{t}\|_{\infty}$. We first bound we bound $\|E^{2}_{t}\|_{\infty}$. By the fact that $\|P(s,a)\|_{1}\leq 1$, $\|V^{(i)}_{t,k}\|_{\infty}$, and $\|V^{\star}\|_{\infty}\leq\frac{1}{1-\gamma}$,
$\|E^{2}_{t}\|_{\infty}$$\leq$ $\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}\|P\|_{1}(\|V^{(i)}_{t,k}\|_{\infty}+\|V^{\star}\|_{\infty})$ (10)
$\leq$ $\frac{2\gamma}{1-\gamma}\sum_{k=0}^{K-1}\eta(1-\eta)^{K-1-k}\leq\frac{2\gamma}{1-\gamma}.$
Next, we bound $\|E^{1}_{t}\|_{\infty}$ by Lemma 4 with $w_{k}^{(i)}=(P_{t,k+1}^{(i)}-P)V^{(i)}_{t,k}$. By the fact that $w_{k}^{(i)}$ is independent of the transition events of other agents, and that $\|P^{(i)}_{t,k}(s,a)\|_{1}\leq 1$, $\|P(s,a)\|_{1}\leq 1$, and $\|V^{(i)}_{t,k}\|_{\infty}\leq\frac{1}{1-\gamma}$, we can prove by proof arguments in Lemma 1 of [68] that $W=\frac{2}{1-\gamma}$ and $\sigma^{2}=\frac{2}{(1-\gamma)^{2}}$. Therefore, for every $s,a,t$
$\|E^{1}_{t}\|_{\infty}$$\leq$ $\gamma\left(\sqrt{\frac{16\eta}{I(1-\gamma)^{2}}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}+\frac{8}{3I(1-\gamma)}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}\right)$ (11)
$\leq$ $\frac{4\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}\right)$
with probability at least $1-\frac{\delta}{|\mathcal{S}||\mathcal{A}|T}$.
Finally, by the union of bounds, we have with probability at least $1-\delta/T$
$\|\Delta_{t,K}\|_{\infty}$$\leq$ $(1-\eta)^{K}\|\Delta_{t,0}\|_{\infty}+\frac{4\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}+\frac{2\gamma}{1-\gamma},$
where $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}$. Since $\Delta_{t,0}=\overline{Q}_{t}-Q^{\star}$, we complete the proof.
Next, Lemma 1 shows that Sparsified-$K$ is UnbiasedComp, whereas Lemma 2 implies that Top-$K$ is BiasedComp. We now provide proof of these lemmas omitted from the main text.
0.A.1.1 Proof of Lemma 1
We obtain the first statement by using (8). Next, we prove the second statement. Since $\mathbf{E}\textsf{Compress}(v)_{j}=v_{j}$ and
$\mathbf{E}|\textsf{Compress}(v)_{j}|^{2}=\frac{1}{p_{i}}v_{i}^{2},$
we have
$\mathbf{E}|\textsf{Compress}(v)_{j}-v_{j}|^{2}$ $=$ $\mathbf{E}|\textsf{Compress}(v)_{j}|^{2}-|v_{j}|^{2}$
$=$ $(1/p_{i}-1)|v_{j}|^{2}$
$\leq$ $(1/p_{\min}-1)|v_{j}|^{2},$
where $p_{\min}=\min_{j=[1,d]}p_{i}$. Finally, we prove the last statement. From the definition of the $\ell_{\infty}$-norm,
$\|\textsf{Compress}(v)-v\|_{\infty}$ $=$ $\max_{j\in[1,d]}|\textsf{Compress}(v)_{j}-v_{j}|$
$\leq$ $\max_{j\in[1,d]}\max(1/p_{j}-1,1)|v_{j}|$
$\leq$ $\max(1/p_{\min}-1,1)\max_{j\in[1,d]}|v_{j}|=q_{\infty}\|v\|_{\infty}.$
0.A.1.2 Proof of Lemma 2
From the definition of the $\ell_{\infty}$-norm and by denoting $I_{K}^{c}(v)=[1,d]-I_{K}(v)$,
$\|\textsf{Compress}(v)-v\|_{\infty}$ $=$ $\max_{j\in I_{K}^{c}(v)}|v_{j}|$
$=$ $\frac{\max_{j\in I_{K}^{c}(v)}|v_{j}|}{\|v\|_{\infty}}\|v\|_{\infty}.$
If $\max_{j\in I_{K}^{c}(v)}|v_{j}|<\|v\|_{\infty}$, then $\|\textsf{Compress}(v)-v\|_{\infty}\leq(1-\alpha)\|v\|_{\infty}$ with $\alpha=1-{\max_{j\in I_{K}^{c}(v)}|v_{j}|}/{\|v\|_{\infty}}$.
To this end, by using the results above, we prove the main theorems in this paper.
0.A.2 Proof of Theorem 5.1
Define $\Delta_{t}=\overline{Q}_{t}-Q^{\star}$. Then, the update for Algorithm 1 with UnbiasedComp can be expressed equivalently as
$\Delta_{t+1}=(1-\beta)\Delta_{t}+\beta\left(\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right)+\beta\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t},$
where $e^{(i)}_{t}=\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t})-(Q_{t,K}^{(i)}-\overline{Q}_{t})$. From the definition of the $\ell_{\infty}$-norm and by the triangle inequality,
$\|\Delta_{t+1}\|_{\infty}\leq(1-\beta)\|\Delta_{t}\|_{\infty}+\beta\left\|\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right\|_{\infty}+\beta\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}.$
To complete the proof, we must bound $\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}$. Since from Lemma 4 of [39], $\|Q^{(i)}_{t,K}-\overline{Q}_{t}\|_{\infty}\leq\|Q_{t,K}^{(i)}\|_{\infty}+\|\overline{Q}_{t}\|_{\infty}\leq\frac{2}{1-\gamma}$. By Lemma 1 we can then prove that
$\mathbf{E}e_{t}^{(i)}$ $=$ 00 0
$\|e_{t}^{(i)}\|_{\infty}$$\leq$ $q_{\infty}\|Q^{(i)}_{t,K}-\overline{Q}_{t}\|_{\infty}\leq\frac{2q_{\infty}}{1-\gamma}$
$\sum_{i=1}^{I}\mathbf{E}(e_{t}^{(i)})^{2}$$\leq$ $\sum_{i=1}^{I}q_{2}\|Q^{(i)}_{t,K}-\overline{Q}_{t}\|_{2}^{2}\leq\frac{4Iq_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}.$
By using Theorem 4 of [68] with $m=1$, we have
$\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}\leq\frac{1}{\sqrt{I}}\left(\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right).$
with probability $1-\delta/(2T)$. Plugging this above result and the result from Lemma 5 with $\delta^{\prime}=\delta/2$ into the main inequality, and using the union of bounds, we have
$\|\Delta_{T}\|_{\infty}$$\leq$ $(1-\beta+\beta(1-\eta)^{K})\|\Delta_{t}\|_{\infty}+\frac{4\beta\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}$
$+\frac{2\beta\gamma}{1-\gamma}+\frac{\beta}{\sqrt{I}}e_{2},$
with probability at least $1-\delta/T$. Here, $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}$ and $e_{2}=\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}$. Finally from Lemma 3, with probability at least $1-\delta$
$\|\Delta_{T}\|_{\infty}$$\leq$ $(1-\beta+\beta(1-\eta)^{K})^{T}\|\Delta_{0}\|_{\infty}+\frac{4\gamma}{C}\sqrt{\frac{\eta}{I}\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}$
$+\frac{2\gamma}{C}+\frac{1}{\sqrt{I}(1-(1-\eta)^{K})}e_{2},$
where $C=(1-\gamma)(1-(1-\eta)^{K})$.
0.A.3 Proof of Theorem 5.2
Define $\hat{\Delta}_{t}=\overline{Q}_{t}+(\beta/I)\sum_{i=1}^{I}e^{(i)}_{t}-Q^{\star}$. Then, the update for Algorithm 1 with BiasedComp can be expressed equivalently as
$\hat{\Delta}_{t+1}=(1-\beta)\hat{\Delta}_{t}+\beta\left(\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right)+\beta^{2}\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t},$
where $e^{(i)}_{t+1}=Q_{t,K}^{(i)}-\overline{Q}_{t}+e^{(i)}_{t}-\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t}+e^{(i)}_{t})$. From the definition of the $\ell_{\infty}$-norm and by the triangle inequality,
$\|\hat{\Delta}_{t+1}\|_{\infty}\leq(1-\beta)\|\hat{\Delta}_{t}\|_{\infty}+\beta\left\|\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right\|_{\infty}+\beta^{2}\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}.$
From Lemma 5, with probability at least $1-\delta/T$
$\|\hat{\Delta}_{t+1}\|_{\infty}$$\leq$ $(1-\beta)\|\hat{\Delta}_{t}\|_{\infty}+\beta(1-\eta)^{K}\|\overline{Q}_{t}-Q^{\star}\|_{\infty}$
$+\frac{4\beta\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}+\frac{2\beta\gamma}{1-\gamma}+\beta^{2}\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty},$
where $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}$. Since
$\|\overline{Q}_{t}-Q^{\star}\|_{\infty}\leq\|\hat{\Delta}_{t}\|_{\infty}+\beta\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty},$
we get: with probability at least $1-\delta/T$
$\|\hat{\Delta}_{t+1}\|_{\infty}$$\leq$ $(1-\beta+\beta(1-\eta)^{K})\|\hat{\Delta}_{t}\|_{\infty}+\frac{4\beta\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}$
$+\frac{2\beta\gamma}{1-\gamma}+\beta^{2}(1+(1-\eta)^{K})\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}.$
Next, suppose $\|\textsf{Compress}(v)-v\|_{\infty}\leq(1-\alpha)\|v\|_{\infty}$ for some $\alpha\in(0,1]$ and $v\in\mathbb{R}^{d}$. Define $w_{t}^{(i)}=e_{t}^{(i)}$. Since from Lemma 4 of [39], $\|Q_{t,K}^{(i)}-\overline{Q}_{t}\|_{\infty}\leq\|Q_{t,K}^{(i)}\|_{\infty}+\|\overline{Q}_{t}\|_{\infty}\leq\frac{2}{1-\gamma}$. Therefore, we have
$\|e^{(i)}_{t+1}\|_{\infty}$$\leq$ $(1-\alpha)\|e^{(i)}_{t}+Q_{t,K}^{(i)}-\overline{Q}_{t}\|_{\infty}$
$\leq$ $(1-\alpha)\|e^{(i)}_{t}\|_{\infty}+(1-\alpha)\|Q_{t,K}^{(i)}-\overline{Q}_{t}\|_{\infty}$
$\leq$ $(1-\alpha)\|e^{(i)}_{t}\|_{\infty}+\frac{2(1-\alpha)}{1-\gamma}.$
From Lemma 3 and by the fact that $e_{0}^{(i)}=0$,
$\|e^{(i)}_{t}\|_{\infty}\leq\frac{2(1-\alpha)}{\alpha(1-\gamma)}.$
Thus,
$\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}$$\leq$ $\frac{1}{I}\sum_{i=1}^{I}\|e^{(i)}_{t}\|_{\infty}\leq\frac{2(1-\alpha)}{\alpha(1-\gamma)}.$
Plugging the upper-bound for $\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}$ into the main inequality yields
$\|\hat{\Delta}_{t+1}\|_{\infty}$$\leq$ $(1-\beta+\beta(1-\eta)^{K})\|\hat{\Delta}_{t}\|_{\infty}+\frac{4\beta\gamma}{1-\gamma}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}$
$+\frac{2\beta\gamma}{1-\gamma}+\beta^{2}(1+(1-\eta)^{K})\frac{2(1-\alpha)}{\alpha(1-\gamma)}$
with probability at least $1-\delta/T$. Using Lemma 3 and the union of bounds, we have the following: with probability at least $1-\delta$
$\|\hat{\Delta}_{T}\|_{\infty}$$\leq$ $(1-\beta+\beta(1-\eta)^{K})^{T}\|\hat{\Delta}_{0}\|_{\infty}$
$+\frac{4}{(1-\gamma)(1-(1-\eta)^{K})}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}$
$+\frac{2\gamma}{(1-\gamma)(1-(1-\eta)^{K})}+\beta(1+(1-\eta)^{K})\frac{2(1-\alpha)}{\alpha(1-\gamma)(1-(1-\eta)^{K})}.$
Since $e^{(i)}_{0}=0$, $\hat{\Delta}_{0}=\Delta_{0}$. Next, by the triangle inequality and by the upper-bound for $\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}$, we have with probability at least $1-\delta$
$\|\Delta_{T}\|_{\infty}$$\leq$ $\|\hat{\Delta}_{T}\|_{\infty}+\frac{\beta}{I}\sum_{i=1}^{I}\|e_{t}^{(i)}\|_{\infty}$
$\leq$ $\|\hat{\Delta}_{T}\|_{\infty}+\frac{2\beta(1-\alpha)}{\alpha(1-\gamma)}$
$\leq$ $(1-\beta+\beta(1-\eta)^{K})^{T}\|\Delta_{0}\|_{\infty}$
$+\frac{4}{(1-\gamma)(1-(1-\eta)^{K})}\sqrt{\frac{\eta}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}$
$+\frac{2\gamma}{(1-\gamma)(1-(1-\eta)^{K})}+\frac{2\beta(1-\alpha)}{\alpha(1-\gamma)}\left(1+\frac{1+(1-\eta)^{K}}{1-(1-\eta)^{K}}\right).$
Appendix 0.B Discussion on the Boundedness of $\|\bar{Q}_{0}\|_{\infty}$
We derive Theorem 5.1 and 5.2 under the assumptions that $\|r\|_{\infty}\leq 1$ and $\|\bar{Q}_{0}\|_{\infty}\leq\frac{1}{1-\gamma}$. These theorems, however, do not lose any generality, since they can be extended to the cases where $0\leq r\leq r_{\max}$ or $|r|\leq r_{\max}$. This can be done by imposing $\|r\|_{\infty}\leq r_{\max}$ and $\|\bar{Q}_{0}\|_{\infty}\leq{r_{\max}}/{(1-\gamma)}$. These results imply that $\|Q^{(i)}_{t,K}-\bar{Q}_{t}\|_{\infty}\leq 2r_{\max}/(1-\gamma)$ and hence that the residual error terms in Theorem 5.1 and 5.2 contain $r_{\max}$ apart from other federated and problem parameters.
Appendix 0.C Details of Experiments
0.C.1 Environments
We use five noisy variants of the grid-world environment, a basic grid-based simulation often used in RL experiments, as shown in Figure 5. In this environment, agents navigate a grid composed of various cells, including walls, empty spaces, and a goal state. Actions available to the agent include moving up, down, left, or right. The state of the environment is represented by the current position of the agent within the grid. Rewards are defined based on the agent’s interactions with the environment; reaching the goal state yields a positive reward $+1$, while taking a non-valid action, such as trying to move through a wall or exit the grid, incurs a penalty $-1$ and the agent stays in the same position. To introduce noise into the environment, a clipped Gaussian noise model is employed, with the specified standard deviation and clipping threshold, set at 0.5 (i.e., $\textsf{clip}\{\mathcal{N}(\mu=0,\sigma^{2}=0.5),-0.5,0.5\}$). This noise affects the agent’s interaction, simulating real-world uncertainties and imperfections in the environment.
(a) Map5×$\times 5$ 5
(b) Map5×$\times 5$ 5w
(c) Map6×$\times 6$ 6w
(d) Map11×$11\times 1$ 11
(e) Map17×$\times 17$ 17w
Figure 5: Various noisy Grid-world environments utilized as experimental tasks, requiring navigation through a grid towards a Goal (green tile) state while avoiding collision with Walls (black tiles) by traversing Empty (white tiles) cells.
0.C.2 Results and Discussion
Applying CompFedRL on our environments has been lead to finding the optimal policy in all cases. For example, see Figure 6 where show $Q$-values in a heatmap with values correspond to colorbar. also arrows showing the action $a$ suggested with the policy learned by our algorithms. In this section, we all have shown the same bath of figures shown in the paper, but for the other four tasks. Results are fully in agreement with what we discussed in Section 6, showing the applicability of algorithms and robustness to the range of task from easy to hard, from loss state-action space to high dimension. To seake of gravity, we omit reporting discussion here and refer the redere to Section 6.
The application of CompFedRL to our task has consistently led to the discovery of optimal policies across all scenarios. For instance, Figure 6 illustrates $Q$-values represented in a heatmap, with corresponding values indicated by the color bar, alongside arrows denoting the suggested action derived from the policy learned by our algorithms, showing the optimality of the founded policy.
This section also presents a cohesive collection of figures akin to those showcased in the paper, each pertaining to one of the remaining four tasks. The outcomes are entirely congruent with the discussions presented in Section 6, underscoring the versatility of the algorithms and their robustness across a spectrum of tasks ranging from straightforward to challenging, and from low-dimensional to high-dimensional state-action spaces. For brevity, we abstain from reiterating the discussions here and direct the reader to Section 6 for further details.
(a) Map5×$\times 5$ 5
(b) Map5×$\times 5$ 5w
(c) Map6×$\times 6$ 6w
(d) Map11×$11\times 1$ 11
(e) Map17×$\times 17$ 17w
Figure 6: Learned $Q$-values and policy for different tasks. Arrows represent the action that learned policy proposes in each state. Here, agents have been trained using CompFedRL algorithm under (a-c )Top-5 (d-e)Top-50  sparsification with the number of agents $I=50$, the number of communication rounds $T=1000$, the number of local epochs $K=10$, the learning rate $\eta=0.01$, and the federated parameter $\beta=0.8$.
0.C.2.1 Results for Map5×$\times 5$ 5 task.
The effects of compression, the number of agents $I$, and the number of local epochs $K$, as well as the impact of the federated parameter $\beta$ and the learning rate $\eta$ on solving the Map5×$\times 5$ 5 task, are depicted in Figures 7, 8, 9, and 10 respectively. It’s noteworthy that due to the smaller state-action spaces in the Map5×$\times 5$ 5 task compared to the Map11×$11\times 1$ 11 task, more compressed operators, specifically Top-5 and Sparsified-5, were utilized instead of Top-50 and Sparsified-50.
(a) Number of local epochs $K=1$
(b) Number of local epochs $K=10$
Figure 7: Impact of compression: the RMSE of the $Q$-estimates with respect to the number of bits communicated per agent for both FedRL (without compression) and CompFedRL under different levels of sparsification via Sparsified-$K$ and Top-$K$ to solve Map5×$\times 5$ 5 task. Here, the number of agents $I=50$, the learning rate $\eta=0.01$, and the federated parameter $\beta=0.8$.
Figure 8: Impact of the number of agents $I$ and the number of local epochs $K$: the RMSE of the $Q$-estimates for CompFedRL under Top-5 sparsification to solve Map5×$\times 5$ 5 task. Here, (a) $T\times K=10000$, the learning rate $\eta=0.01$, and the federated parameter $\beta=1$. (b) the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
(a) Top-5
(b) Sparsified-5
Figure 9: Impact of federated parameter $\beta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-5  and Sparsified-5  sparsification to solve Map5×$\times 5$ 5 task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the learning rate $\eta=0.1$.
(a) Top-5
(b) Sparsified-5
Figure 10: Impact of learning rate $\eta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-5  and Sparsified-5  sparsification to solve Map5×$\times 5$ 5 task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
0.C.2.2 Results for Map5×$\times 5$ 5w task.
The impacts of compression, the number of agents $I$, and the number of local epochs $K$, alongside the influence of the federated parameter $\beta$ and the learning rate $\eta$ on addressing the Map5×$\times 5$ 5w task, are illustrated in Figures 11, 12, 13, and 14, respectively. It is worth noting that due to the smaller state-action spaces in the Map5×$\times 5$ 5w task compared to the Map11×$11\times 1$ 11 task, more compressed operators—specifically Top-5 and Sparsified-5—were employed instead of Top-50 and Sparsified-50, respectively.
(a) Number of local epochs $K=1$
(b) Number of local epochs $K=10$
Figure 11: Impact of compression: the RMSE of the $Q$-estimates with respect to the number of bits communicated per agent for both FedRL (without compression) and CompFedRL under different levels of sparsification via Sparsified-$K$ and Top-$K$ to solve Map5×$\times 5$ 5w task. Here, the number of agents $I=50$, the learning rate $\eta=0.01$, and the federated parameter $\beta=0.8$.
Figure 12: Impact of the number of agents $I$ and the number of local epochs $K$: the RMSE of the $Q$-estimates for CompFedRL under Top-5 sparsification to solve Map5×$\times 5$ 5w task. Here, (a) $T\times K=10000$, the learning rate $\eta=0.01$, and the federated parameter $\beta=1$. (b) the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
(a) Top-5
(b) Sparsified-5
Figure 13: Impact of federated parameter $\beta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-5  and Sparsified-5  sparsification to solve Map5×$\times 5$ 5w task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the learning rate $\eta=0.1$.
(a) Top-5
(b) Sparsified-5
Figure 14: Impact of learning rate $\eta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-5  and Sparsified-5  sparsification to solve Map5×$\times 5$ 5w task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
0.C.2.3 Results for Map6×$\times 6$ 6w task.
The impacts of compression, the number of agents $I$, and the number of local epochs $K$, along with the influence of the federated parameter $\beta$ and the learning rate $\eta$ on addressing the Map6×$\times 6$ 6w task, are illustrated in Figures 15, 16, 17, and 18 respectively. Notably, due to the smaller state-action spaces in the Map6×$\times 6$ 6w task compared to the Map11×$11\times 1$ 11 task, more compressed operators, specifically Top-5 and Sparsified-5, were employed instead of Top-50 and Sparsified-50.
(a) Number of local epochs $K=1$
(b) Number of local epochs $K=10$
Figure 15: Impact of compression: the RMSE of the $Q$-estimates with respect to the number of bits communicated per agent for both FedRL (without compression) and CompFedRL under different levels of sparsification via Sparsified-$K$ and Top-$K$ to solve Map6×$\times 6$ 6w task. Here, the number of agents $I=50$, the learning rate $\eta=0.01$, and the federated parameter $\beta=0.8$.
Figure 16: Impact of the number of agents $I$ and the number of local epochs $K$: the RMSE of the $Q$-estimates for CompFedRL under Top-5 sparsification to solve Map6×$\times 6$ 6w task. Here, (a) $T\times K=10000$, the learning rate $\eta=0.01$, and the federated parameter $\beta=1$. (b) the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
(a) Top-5
(b) Sparsified-5
Figure 17: Impact of federated parameter $\beta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-5  and Sparsified-5  sparsification to solve Map6×$\times 6$ 6w task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the learning rate $\eta=0.1$.
(a) Top-5
(b) Sparsified-5
Figure 18: Impact of learning rate $\eta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-5  and Sparsified-5  sparsification to solve Map6×$\times 6$ 6w task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
0.C.2.4 Results for Map17×$\times 17$ 17w task
The impacts of compression, the number of agents $I$, and the number of local epochs $K$, along with the influence of the federated parameter $\beta$ and the learning rate $\eta$ on solving the Map17×$\times 17$ 17w task, are illustrated in Figures 19, 20, 21, and 22 respectively.
(a) Number of local epochs $K=1$
(b) Number of local epochs $K=10$
Figure 19: Impact of compression: the RMSE of the $Q$-estimates with respect to the number of bits communicated per agent for both FedRL (without compression) and CompFedRL under different levels of sparsification via Sparsified-$K$ and Top-$K$ to solve Map17×$\times 17$ 17w task. Here, the number of agents $I=50$, the learning rate $\eta=0.01$, and the federated parameter $\beta=0.8$.
Figure 20: Impact of the number of agents $I$ and the number of local epochs $K$: the RMSE of the $Q$-estimates for CompFedRL under Top-50 sparsification to solve Map17×$\times 17$ 17w task. Here, (a) $T\times K=10000$, the learning rate $\eta=0.01$, and the federated parameter $\beta=1$. (b) the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.
(a) Top-50
(b) Sparsified-50
Figure 21: Impact of federated parameter $\beta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-50  and Sparsified-50  sparsification to solve Map17×$\times 17$ 17w task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the learning rate $\eta=0.1$.
(a) Top-50
(b) Sparsified-50
Figure 22: Impact of learning rate $\eta$: the RMSE of the $Q$-estimates with respect to the communication round for CompFedRL under Top-50  and Sparsified-50  sparsification to solve Map17×$\times 17$ 17w task. Here, the number of agents $I=50$, the number of communication rounds $T=10000$, the number of local epochs $K=1$, and the federated parameter $\beta=0.8$.